OpenStudy (anonymous):

cot ^-1(square root of 3)

OpenStudy (anonymous):

find a number whose tangent is \(\frac{1}{\sqrt3}\)

OpenStudy (anonymous):

look on the unit circle to see where \[\sin(x)=\frac{1}{2},\cos(x)=\frac{\sqrt3}{2}\] that is the number you are looking for

OpenStudy (anonymous):

yeah I know that but im doing the hw online and it keep saying that im wrong

OpenStudy (anonymous):

did you put in \(\frac{\pi}{6}\)?

OpenStudy (anonymous):

yes, i got it now. Thank you!

OpenStudy (anonymous):

yw